### v2

# Cooperative Nonlinear Guidance Strategies for Guaranteed Pursuit-Evasion

Abstract
This paper addresses the pursuit-evasion problem involving three agents– a purser, an evader, and a defender. We develop cooperative guidance laws for the evader-defender team that guarantee that the defender intercepts the pursuer before it reaches the vicinity of the evader. Unlike heuristic methods, optimal control, differential game formulation, and recently proposed time-constrained guidance techniques, we propose a geometric solution to safeguard the evader from the pursuer’s incoming threat. The proposed strategy is computationally efficient and expected to be scalable as the number of agents increases. Another alluring feature of the proposed strategy is that the evader-defender team does not require the knowledge of the pursuer’s strategy and that the pursuer’s interception is guaranteed from arbitrary initial engagement geometries. We further show that the necessary error variables for the evader-defender team vanish within a time that can be exactly prescribed prior to the three-body engagement. Finally, we demonstrate the efficacy of the proposed cooperative defense strategy via simulation in diverse engagement scenarios.
Index Terms:
Pursuit-evasion, autonomy, aerospace, multiagent systems, aircraft defense.
I Introduction
The pursuit-evasion problem has long been a subject of interest in game theory, robotics, aerospace, and control engineering. While the pursuit-evasion scenarios involving only two participants have been extensively explored, the extension to a three-body pursuit-evasion problem adds a layer of complexity and requires higher autonomy and strategic decision-making. Such engagements involve three agents, namely, a pursuer, a target, and a defender. The pursuer tries to capture the target, which in turn is assisted by a defender to avoid getting captured.
Early research on the kinematics of three-body engagement could be found in [1]. In [1], a closed-form solution was derived for constant-bearing collision courses, while the work in [2] focused on determining the intercept point’s location in the evader-centered reference frame. For three-agent engagements, optimal control-based formulations with specific objectives, such as minimizing energy or cost, were employed in cooperative guidance strategies, as discussed in [3, 4, 5, 6, 7]. In [3], the authors presented a cooperative optimal guidance strategy integrated with a differential game formulation to maximize the separation between pursuer and evader. In [4], optimal cooperative pursuit-evasion strategies for the defender-evader team were proposed, considering arbitrary order-linearized dynamics for each agent. It was assumed that the pursuer’s guidance strategy was known in this case. The work in [5] introduced a multiple-model adaptive estimator approach for cooperative information-sharing between the evader and defender to estimate the likely linear guidance strategy of the pursuer. The work in [6] discussed three-layer cooperation between the defender and evader and explored information exchange between them, whereas that in [7] provided algebraic conditions under which the pursuer could capture the evader by bypassing the defender. Note that most of these strategies relied on linearized dynamics, simplifying guidance design but potentially limiting their applicability in diverse operating conditions and scenarios with significant heading errors.
Guidance strategies developed in a nonlinear context can overcome these limitations and enhance performance, e.g., by relaxing small heading error assumptions and accounting for turn constraints. Notable works in this regard include [8, 9, 10, 11, 12]. In [8], the authors introduced a sliding mode control-based terminal intercept guidance and autopilot design for defenders to protect the evader from incoming pursuers. Another nonlinear guidance strategy employing sliding mode control was discussed in [9]. In [10], a nonlinear guidance strategy was explored for scenarios where multiple defenders simultaneously intercept the pursuer before it reaches the evader. In [11], nonlinear feedback laws were developed to guide the evader on a collision course with the pursuer as a decoy, allowing the defender to intercept the pursuer before it captures the evader. This approach also provided the defender with the flexibility to adopt either a defensive or aggressive stance based on mission requirements. Another nonlinear guidance strategy, based on relative line-of-sight error and time-to-go deviation between the evader and the defender, was presented in [12]. It is important to note that these guidance strategies relied on time-to-estimates, which may not always be available with the required precision, potentially affecting their effectiveness.
Geometrical approaches have also found application in three-agent pursuit-evasion scenarios. For instance, in [13, 14], a method centered on line-of-sight guidance, a three-point guidance strategy, was explored. These three points were defined as the evader, the defender, and the pursuer. The approach demonstrated that if the defender remains aligned with the line-of-sight connecting the evader and the pursuer, the interception of the pursuer is assured before it nears the evader. In a related vein, a modified version of the command-to-line-of-sight guidance approach, incorporating optimal control theory and velocity error feedback, was introduced in [15]. Furthermore, in [16], the authors presented a guidance strategy based on a barrier Lyapunov function to protect the evader from a pursuer. It is worth noting that the use of the barrier function imposed restrictions on some engagement variables, such as the defender’s initial heading angle, thereby potentially limiting the target set in which the game could terminate. Motivated by these results, the focus of the current paper is analyzing and presenting a simple and intuitive geometry-based solution for guaranteed pursuit evasion. The merits of this work can be succinctly summarized as follows:
  • We propose a geometrical approach to guarantee pursuit-evasion from arbitrary three-body engagement geometries. The proposed solution, which is the evader-defender cooperative guidance strategy, ensures that the defender always arrives at a certain angle within a prescribed time, regardless of the initial geometry, thereby preventing the pursuer from capturing the evader.
  • Unlike LOS angle-based geometric guidance, wherein the defender has to strictly maintain a fixed angle of $\pi$ with respect to the pursuer-evader LOS, the proposed strategy is less stringent and only requires the said angle to be within a broad interval of $\left[{\pi}/{2},{3\pi}/{2}\right]$. This allows the defender to have more flexibility in desired angle selection depending on engagement scenarios.
  • Within our problem framework, the dynamics governing the agents are inherently nonlinear and account for large heading angle errors and non-holonomic constraints. Consequently, the steering control variable for each agent is its lateral acceleration, a pragmatic choice when compared to the manipulation of heading angles. Such consideration is more practical in the context of aerial vehicles, e.g., in aircraft defense.
  • The proposed strategy has an intuitive appeal, is expected to be computationally efficient and sets itself apart from heuristic methodologies, optimal control strategies, and formulations rooted in differential games, where analytical solutions may cease to exist due to challenges associated with nonlinearity and the complexities inherent in non-convex optimization.
  • The proposed geometry-based solutions are versatile and can be applied to a wide range of pursuit-evasion scenarios involving different numbers of agents, dimensions, and constraints. By analyzing the geometry of the problem, the results in this paper open up new avenues to identify necessary and sufficient conditions or configurations that lead to successful evasion or capture.
Note that even if the pursuer is sufficiently close to the evader, the evader and defender will cooperatively maneuver such that the defender can intercept the pursuer. Of course, there could be a few cases where the pursuer is too close to the evader, while the defender is too far away from them. In such cases, it is apparent that the defender may not be able to defend the evader. However, it is also worthwhile to note that such cases are impractical to consider from a three-body aircraft defense scenario because if the defender is too far away, then the engagement, for all practical purposes, is only for two agents.
II Problem Formulation
We consider a cooperative defense problem involving nonholonomic agents, namely, a pursuer (P), an evader (E), and a defender (D). The pursuer aims to intercept the evader. In contrast, the defender’s objective is to neutralize the pursuer before it reaches the vicinity of the evader. Thus, the evader and the defender are allies, whereas the pursuer is the opposition. Such a scenario leads to an planar (2D) engagement, as shown in Figure 1. The agents evolve according to
$\dot{x}_{i}=v_{i}\cos\gamma_{i},~{}\dot{y}_{i}=v_{i}\sin\gamma_{i},~{}\dot{\gamma}_{i}=\dfrac{a_{i}}{v_{i}};\;\forall i={\mathrm{P},\mathrm{E},\mathrm{D}},$ (1)
where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$, $v_{i}\in\mathbb{R}_{+}$, and $\gamma_{i}\in(-\pi,\pi]$ denote the position, speed, and the heading angle of the $i$th agent, respectively, whereas $a_{i}$ is its steering control (lateral acceleration), which is assumed to be bounded. Thus, $\left|a_{i}\right|\leq a_{i}^{\max}\in\mathbb{R}_{+}$. This is unlike previous studies where the agents have simple dynamics, and their instantaneous heading angles were used to control them [17, 18].
Figure 1: Three-body engagement.
Note that the considerations herein lead to nonlinear dynamics, including the vehicles’ turning constraints, and finding analytical solutions to guaranteed winning zones for either party is generally NP-hard. In this paper, the speeds and maneuverability of the agents are such that $v_{\mathrm{P}}\approx v_{\mathrm{D}}>v_{\mathrm{E}}$, and $a_{\mathrm{P}}^{\max}\approx a_{\mathrm{D}}^{\max}>a_{\mathrm{E}}^{\max}$, essentially implying that the pursuer and the defender are similar in capabilities with a speed advantage over the evader. As seen from Figure 1, the agents have relative separations, $r_{\ell}$, and the line-of-sight (LOS) angles between any two pair of agents is $\lambda_{\ell}$, where $\ell=\mathrm{EP},\mathrm{DE},$ and $\mathrm{DP}$. The engagement kinematics governing the relative motion between any two pairs can be expressed in polar coordinates as
$\dot{r}_{\mathrm{EP}}$ $=v_{\mathrm{P}}\cos\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)-v_{\mathrm{E}}\cos\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}\right)$
$=v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}-v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}=v_{r_{\mathrm{EP}}},$ (2a)
$r_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}$ $=v_{\mathrm{P}}\sin\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)-v_{\mathrm{E}}\sin\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}\right)$
$=v_{\mathrm{P}}\sin\delta_{\mathrm{PE}}-v_{\mathrm{E}}\sin\delta_{\mathrm{EP}}=v_{\lambda_{\mathrm{EP}}},$ (2b)
$\dot{r}_{\mathrm{DP}}$ $=v_{\mathrm{P}}\cos\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}\right)-v_{\mathrm{D}}\cos\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}\right)$
$=v_{\mathrm{P}}\cos\delta_{\mathrm{PD}}-v_{\mathrm{D}}\cos\delta_{\mathrm{DP}}=v_{r_{\mathrm{DP}}},$ (2c)
$r_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}$ $=v_{\mathrm{P}}\sin\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}\right)-v_{\mathrm{D}}\sin\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}\right)$
$=v_{\mathrm{P}}\sin\delta_{\mathrm{PD}}-v_{\mathrm{D}}\sin\delta_{\mathrm{DP}}=v_{\lambda_{\mathrm{DP}}},$ (2d)
$\dot{r}_{\mathrm{DE}}$ $=v_{\mathrm{E}}\cos\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}\right)-v_{\mathrm{D}}\cos\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}\right)$
$=v_{\mathrm{E}}\cos\delta_{\mathrm{ED}}-v_{\mathrm{D}}\cos\delta_{\mathrm{DE}}=v_{r_{\mathrm{DE}}},$ (2e)
$r_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}$ $=v_{\mathrm{E}}\sin\left(\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}\right)-v_{\mathrm{D}}\sin\left(\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}\right)$
$=v_{\mathrm{E}}\sin\delta_{\mathrm{ED}}-v_{\mathrm{D}}\sin\delta_{\mathrm{DE}}=v_{\lambda_{\mathrm{DE}}},$ (2f)
where $v_{r_{\mathrm{EP}}},v_{r_{\mathrm{DP}}},v_{r_{\mathrm{DE}}},v_{\lambda_{\mathrm{EP}}},v_{\lambda_{\mathrm{DP}}}$, and $v_{\lambda_{\mathrm{DE}}}$, represent the components of relative velocities of the relevant agents along and perpendicular to the corresponding LOS of their respective engagements. The quantities $\delta_{k}$ in (2) denote the corresponding lead angles and are defined as $\delta_{\mathrm{PE}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{EP}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{PD}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{DP}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{ED}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}$, $\delta_{\mathrm{DE}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}.$ Note that (2a)–(2b) describe the equations of motion between the evader and the pursuer, while (2c)–(2d) represent the same for the defender-pursuer pair. On the other hand, the cooperative engagement between the evader-defender pair is described by (2e)–(2f).
The goal of this paper is to find a feasible nonlinear guidance strategy such that the evader-defender team could cooperatively ensure the evader’s survival regardless of the pursuer’s guidance law. This essentially means that we are interested in designing $a_{\mathrm{E}}$ and $a_{\mathrm{D}}$ such that a target set $\mathscr{T}=\{\mathscr{S}\,|\,r_{\mathrm{DP}}(t_{f})=0\}$, where $\mathcal{S}$ is the set of relevant states of the agents (e.g., position, velocity, heading, range, LOS, etc.) and $t_{f}$ is the time when the defender captures the pursuer, can be reached. Note that the evader-defender alliance has no knowledge of the pursuer’s guidance law. However, each agent can measure the relative information of every other agent.
III Main Results
Consider the triangle formed by joining the pursuer, the evader, and the defender at any given point of time, as illustrated in $\triangle$ P1E1D1 in Figure 2. A circle can be thought of as circumscribing this triangle such that the relative distances between the pair of agents form chords in the circle. The relative distance between the pursuer and the evader, $r_{\mathrm{E}_{1}\mathrm{P}_{1}}$, can be considered the base of $\triangle$ P1E1D1, whereas the other two relative distances represent the other sides of the triangle. It is immediate from basic geometry that when a chord divides a circle, it creates two segments– one with an acute angle subtended (major segment) and the other with an obtuse angle (minor segment). The diameter is the longest chord, and it subtends a right angle in both segments. Following this rule, it is apparent that $\angle$ E1D1P1 is obtuse, whereas $\angle$ E0D0P0 in $\triangle$ P0E0D0 subtends $90^{\circ}$ as E0P0 passes through the diameter of the circle.
Let us define an angle $\chi$ such that it is the angle subtended by the evader-defender and the defender-pursuer LOS at any segment, which is measured positive in the counterclockwise sense. Referring to Figure 2, it is the $\angle$ E1D1P1 in $\triangle$ P1E1D1 and $\angle$ E0D0P0$=90^{\circ}$ in $\triangle$ P0E0D0. In the case of $\triangle$ P2E2D2, the fact that interior $\angle$ E2D2P2 is obtuse still holds but $\chi$ is exterior of this angle to respect the consistency of our definition. The following proposition provides a sufficient condition for the defender to capture the pursuer before the latter could intercept the evader.
Figure 2: Illustration of the angle $\chi$.
Proposition 1.
Consider the three-body engagement described using (2). If the defender maintains a positive closing speed with respect to the pursuer and attains an angle $\chi\in\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$, then the pursuer’s capture is guaranteed before the evader could be captured.
Proof.
Referring to Figure 2, it is immediate that if $\chi\in\left(\dfrac{\pi}{2},{\pi}\right)$ represents a scenario depicted in $\triangle$ P1E1D1. It follows that the side E1P1 is the longest. Hence, E1P1$>$D1P1. If the evader-defender team cooperates such that the defender maintains a fixed $\chi\in\left(\dfrac{\pi}{2},{\pi}\right)$ while also ensuring a positive closing speed with respect to the pursuer ($v_{r_{\mathrm{DP}}}<0$), the $\triangle$ P1E1D1 will shrink in proportion and may change its orientation, thereby generating smaller similar triangles. Eventually, D1P1 will degenerate to zero (or $r_{\mathrm{D}_{1}\mathrm{P}_{1}}\to 0$) while E1P1 (or $r_{\mathrm{E}_{1}\mathrm{P}_{1}}$) will still be positive. By a similar argument, it readily follows that $r_{\mathrm{D}_{2}\mathrm{P}_{2}}\to 0$ before $r_{\mathrm{E}_{2}\mathrm{P}_{2}}$ if $\chi\in\left({\pi},\frac{3\pi}{2}\right)$. The cases of $\chi=\frac{\pi}{2},\pi$ are extremes but no different. When $\chi=\dfrac{\pi}{2}$ (as in $\triangle$ P0E0D0), the side E0P0 is the hypotenuse and still the longest, so same reasoning can be applied. Finally, when $\chi=\pi$, the defender is always directly between and on the pursuer-evader LOS, resulting in the pursuer’s capture by the defender prior to the interception of the evader by the pursuer. ∎
From Figure 1, it is imperative to mathematically define the angle $\chi$ as
$\chi=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{DE}}$ (3)
for a general case. Toward this end, our goal is to ensure that the defender always attains a fixed angle, say, $\chi^{\star}\in\left[\frac{\pi}{2},\frac{3\pi}{2}\right]$. Thus, the control objective is to nullify the error variable,
$\beta=\chi-\chi^{\star}=\pi+\lambda_{\mathrm{DP}}-\lambda_{\mathrm{DE}}-\chi^{\star}.$ (4)
Additionally, we also desire that $\beta\to 0$ within a time that is independent and uniform with respect to the initial three-body engagement geometry for a guaranteed pursuit-evasion.
Lemma 1.
The dynamics of the LOS angle in each pair of engagements have a relative degree of two with respect to the relevant steering controls in that engagement.
Proof.
On differentiating the LOS rate of evader-pursuer engagement given in (2b) with respect to time, one may obtain
$r_{\mathrm{EP}}\ddot{\lambda}_{\mathrm{EP}}+\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}=v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}\dot{\delta}_{\mathrm{PE}}-v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}\dot{\delta}_{\mathrm{EP}}.$ (5)
Using the fact that $\dot{\delta}_{\mathrm{PE}}=\dot{\gamma}_{\mathrm{P}}-\dot{\lambda}_{\mathrm{EP}}$ and $\dot{\delta}_{\mathrm{EP}}=\dot{\gamma}_{\mathrm{E}}-\dot{\lambda}_{\mathrm{EP}}$, together with (1), the expression in (5) becomes
$r_{\mathrm{EP}}\ddot{\lambda}_{\mathrm{EP}}=$ $~{}-\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}+v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}\left(\dfrac{a_{\mathrm{P}}}{v_{\mathrm{P}}}-\dot{\lambda}_{\mathrm{EP}}\right)$
$~{}-v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}\left(\dfrac{a_{\mathrm{E}}}{v_{\mathrm{E}}}-\dot{\lambda}_{\mathrm{EP}}\right).$ (6)
After arranging the similar terms together in (6), we get
$\ddot{\lambda}_{\mathrm{EP}}=$ $~{}\dfrac{\dot{\lambda}_{\mathrm{EP}}}{r_{\mathrm{EP}}}\left(-\dot{r}_{\mathrm{EP}}-v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}+v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}\right)$
$~{}+\dfrac{\cos\delta_{\mathrm{PE}}}{r_{\mathrm{EP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{EP}}}{r_{\mathrm{EP}}}a_{\mathrm{E}},$ (7)
which can be simplified using (2a) to
$\ddot{\lambda}_{\mathrm{EP}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}}{r_{\mathrm{EP}}}+\dfrac{\cos\delta_{\mathrm{PE}}}{r_{\mathrm{EP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{EP}}}{r_{\mathrm{EP}}}a_{\mathrm{E}}.$ (8)
A similar procedure results in
$\ddot{\lambda}_{\mathrm{DP}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}a_{\mathrm{D}},$ (9)
$\ddot{\lambda}_{\mathrm{DE}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}-\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}a_{\mathrm{D}}.$ (10)
One may notice from (8)–(10) that the dynamics of LOS angles have a relative degree of two with respect to the lateral acceleration (steering controls) of the corresponding agents. This concludes the proof. ∎
Using the results in Lemma 1, we can now obtain the dynamics of the error variable, $\beta$, as
$\ddot{\beta}=$ $~{}\ddot{\lambda}_{\mathrm{DP}}-\ddot{\lambda}_{\mathrm{DE}}$ (11)
$=$ $~{}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}$
$~{}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}},$ (12)
which is affine with the steering controls of the evader and the defender, whereas that of the pursuer is an unknown quantity. The dynamics of $\beta$ can be written as
$\ddot{\beta}=$ $~{}\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}+\mathcal{U},$ (13)
where $\mathcal{U}$ is defined as the joint cooperative maneuver (or net control effort) given by
$\mathcal{U}=-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}.$ (14)
In order to nullify $\beta$, we need to design $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$ such that the evader-defender team cooperatively maneuvers to place the defender on a fixed $\chi^{\star}$ in the interval $\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$ regardless of the initial three-body engagement geometry.
Toward this objective, we consider a dual-layer sliding manifold, with the inner-layer as
$\underaccent{\bar}{\mathcal{S}}(t)=\dot{\beta}(t)-g(t)$ (15)
where $\dot{\beta}(t)=\dot{\lambda}_{\mathrm{DP}}-\dot{\lambda}_{\mathrm{DE}}$ and $g(t)$ is given by
$g(t)=\begin{cases}-\dfrac{k_{1}\,\beta(t)}{t^{\star}-t};&0\leq t<t^{\star},\\ \phantom{t-t}0;&t\geq t^{\star},\end{cases}$ (16)
for some $k_{1}\in\mathbb{N}$. Note that the variable $t^{\star}$ denotes the time instant at which the error variable $\beta(t)$ nullifies to zero. As a consequence, the defender attains a predefined angle $\chi^{\star}\in\left[\pi/2,3\pi/2\right]$ within the same time $t^{\star}$. Thereafter, to eliminate the reaching phase and achieve a global exact-time convergence, we consider the outer-layer sliding manifold as
$\bar{\mathcal{S}}(t)=\underaccent{\bar}{\mathcal{S}}(t)+h(t),$ (17)
where $h(t)$ is an auxiliary function defined as
$h(t)=\begin{cases}\dfrac{h(0)}{t_{1}^{k_{2}}}\left(t_{1}-t\right)^{k_{2}};&0\leq t<t_{1},\\ \phantom{t-t}0;&t\geq t_{1},\end{cases}$ (18)
for some $k_{2}\in\mathbb{N}$, and $h(0)$ is designed to ensure $\bar{\mathcal{S}}(0)=0$, which can be obtained from (17) as $h(0)=-\underaccent{\bar}{\mathcal{S}}(0)$. In (18), $t_{1}$ is the exact time instant from which onward the inner layer sliding manifolds become identically equal to zero. In other words, the sliding mode is enforced on the inner-layer sliding manifold at time $t=t_{1}$. The fundamental concept behind eliminating the reaching phase is to ensure global exact-time convergence for the sliding variables and achieve a complete sliding mode response. This, in consequence, bolsters robustness, as sliding mode controllers typically experience diminished robustness during their reaching phase. We are now ready to present the joint cooperative effort of the evader-defender alliance next.
Theorem 2.
Consider the three-body engagement described using (2), the dynamics of the error variable, (13), and the dual-layer sliding manifold in (15) and (17). If the joint cooperative maneuver of the evader-defender team is designed as
$\mathcal{U}=-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}-\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dot{g}(t)-\dot{h}(t)-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}}),$ (19)
for some $\mathcal{K}>\sup_{t\geq 0}\left({a_{\mathrm{P}}^{\max}}/{r_{\mathrm{DP}}}\right)$, then the evader-defender team cooperatively maneuvers such that the defender converges to the desired angle $\chi^{\star}\in\left[{\pi}/{2},{3\pi}/{2}\right]$, within a time $t^{\star}$ prescribed prior to the engagement, regardless of the three-body initial engagement geometry.
Proof.
Consider a continuous, radially unbounded Lyapunov function candidate $V=|\bar{\mathcal{S}}|$. Hereafter, we streamline the notation again by dropping the arguments of variables denoting their time dependency. Upon differentiating $V$ with respect to time and using the relations in (15) and (17), one may obtain
$\dot{V}=\mathrm{sign}(\bar{\mathcal{S}})\dot{\bar{\mathcal{S}}}=\mathrm{sign}(\bar{\mathcal{S}})\left(\dot{\underaccent{\bar}{\mathcal{S}}}+\dot{h}\right)=\mathrm{sign}(\bar{\mathcal{S}})\left(\ddot{\beta}-\dot{g}+\dot{h}\right),$ (20)
which can be further written using the expression in (13) as
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}\right.$
$~{}\left.+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dot{g}+\dot{h}\right].$ (21)
Upon substituting the expression of $\mathcal{U}$ from (19) in the above relation, the derivative of the Lyapunov function candidate reduces to
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}\right]$
$\leq$ $-\mathcal{K}+\dfrac{1}{r_{\mathrm{DP}}}a_{\mathrm{P}}^{\max}~{}<0,$ (22)
[…]
Now, for a time instant $t=t_{1}<t^{\star}$, the reduced order dynamics from (17) can be expressed as
$\dot{\beta}-g=0\implies\dot{\beta}+\dfrac{k_{1}\,\beta}{t^{\star}-t}=0.$ (23)
On integrating (23) within suitable limits, one may obtain
$\int_{\beta(0)}^{\beta(t)}\dfrac{\dot{\beta}}{\beta}\mathrm{d}t=-\int_{0}^{t}\dfrac{k_{1}}{t^{\star}-t}\mathrm{d}t,$ (24)
which results in
$\ln\left[\dfrac{\beta(t)}{\beta(0)}\right]=\ln\left[\dfrac{t^{\star}-t}{t^{\star}}\right]^{k_{1}}.$ (25)
After some simplifications, one can express
$\beta(t)=\dfrac{\beta(0)}{t^{\star^{k_{1}}}}\left(t^{\star}-t\right)^{k_{1}},$ (26)
which clearly shows that at time instant $t=t^{\star}$, $\beta(t)=0$ regardless of $\beta(0)$. Thus, the defender attains $\chi^{\star}$ within $t^{\star}$ due to the joint cooperative maneuver of the evader-defender team. This concludes the proof. ∎
Note that $\mathcal{U}$ in Theorem 2 is the net control authority that the evader-defender pair has to maintain in order to achieve a successful interception of the pursuer. Next, we endeavor to allocate this net control effort to the evader and the defender. Needless to say, there can be several choices for the allocation of the total control effort between the defender and the evader. However, we focus on a control allocation technique that instantaneously minimizes the $\mathcal{L}_{2}$ norm of the net control subject to the affine constraint (14). Towards that, our aim is to minimize
$\mathcal{C}:=\sqrt{a_{\mathrm{E}}^{2}+a_{\mathrm{D}}^{2}}.$ (27)
The essence of this effort allocation is presented in the next theorem.
Theorem 3.
The evader-defender team cooperatively ensures the interception of the pursuer with minimum values of the lateral accelerations,
$a_{\mathrm{E}}=$ $\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]}$
$+\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]}$
$-\dfrac{r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{g}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{h}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})}{\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]},$ (28)
$a_{\mathrm{D}}=$ $\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}$
$+\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}$
$-\dfrac{r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{g}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{h}(t)-\mathcal{K}r_{\mathrm{DE}}r_{\mathrm{DP}}\mathrm{sign}(\bar{\mathcal{S}})}{\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}.$ (29)
Proof.
From (14), one may write
$a_{\mathrm{E}}=\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}\left[\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}-\mathcal{U}\right].$ (30)
Now substituting the above expression of $a_{\mathrm{E}}$ in the cost function (27) yields which can also be written as $\mathcal{C}=\sqrt{\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)^{2}+a_{\mathrm{D}}^{2}}$, where $\mathfrak{B}_{1}=\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}$ and $\mathfrak{B}_{2}=\dfrac{r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DE}}r_{\mathrm{DP}}}$. Computing the partial derivative of $\mathcal{C}$ with respect to $a_{\mathrm{D}}$ and equating it to zero leads to
$\dfrac{\partial\mathcal{C}}{\partial a_{\mathrm{D}}}=\dfrac{1}{2\mathcal{C}}\left[2\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)\mathfrak{B}_{2}+2a_{\mathrm{D}}\right]=0$
$\implies\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}a_{\mathrm{D}}+a_{\mathrm{D}}=\mathfrak{B}_{1}\mathcal{U}\implies a_{\mathrm{D}}=\dfrac{\mathfrak{B}_{1}\mathcal{U}}{1+\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}},$ (31)
which on substituting the value of $\mathcal{U}$ from (19) and $\mathfrak{B}_{1}$ and $\mathfrak{B}_{2}$ leads to the expression of the defender’s lateral acceleration in (29). Thereafter, one may use (30) to arrive at (28). It can be readily verified using higher-order derivative tests that the values thus obtained are indeed minima. By following a similar procedure, one can obtain the value of $a_{\mathrm{E}}$. This concludes the proof. ∎
IV Simulations
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls $a_{\rm D}$ and $a_{\rm E}$.
(d) $\mathcal{U}$ and $a_{\rm P}$.
Figure 3: The defender intercepts the pursuer at various values of the angle $\chi^{\star}$.
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls of the evader and the defender.
(d) Joint effort of the evader-defender team and the pursuer’s maneuver.
Figure 4: The pursuer uses different guidance strategies.
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls of the defender and the pursuer.
Figure 5: The defender safeguards the stationary evader from different initial configurations.
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls of the defender and the pursuer.
Figure 6: The defender safeguards the stationary evader from the same initial configuration.
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls $a_{\rm D}$ and $a_{\rm E}$.
(d) $\mathcal{U}$ and $a_{\rm P}$.
Figure 7: The defender intercepts the pursuer with reduced maximum acceleration
We now demonstrate the performance of the proposed nonlinear guidance strategy using simulation for various three-body engagement scenarios. We assume $v_{\mathrm{P}}=v_{\mathrm{D}}=200$ m/s, whereas $v_{\mathrm{E}}=100$ m/s. In each case, we place the evader at the origin of the inertial place at the beginning of the engagement. The initial positions of the agents in the trajectory plots are denoted using square markers, whereas an intercept is represented using an asterisk. In the following plots, the pursuer’s trajectories and inputs are shown in red, while those of the evader and defender are shown in blue and green, respectively. In our simulations, the pursuer and the defender possess the same maneuverability capabilities, limited to ±$\pm 20$ g. On the other hand, the evader can only apply a maximum steering control of $10$ g in either direction. Here g$=9.81$ m/s2 is the acceleration due to gravity. Thus, the pursuer and the defender are evenly matched in terms of both speed and acceleration capabilities. The design parameters are chosen to be $t_{1}=3$ s, $t^{\star}=6$ s, $k_{1}=6$, $k_{2}=3$ and $\mathcal{K}=5$. Such choices of $t_{1}$ and $t^{\star}$ ensure that the sliding manifold converges to zero within 3 s, whereas the error vanishes within 6 s.
Consider a scenario where the pursuer is using a guidance strategy that is a function of the pursuer-evader LOS rate ($\dot{\lambda}_{\mathrm{EP}}$), for example, proportional-navigation guidance. For the pursuer, such a strategy may also be optimal in certain cases as it aims to arrive on the collision course with minimum effort. The guidance command for the pursuer is given by $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, where $\mathscr{N}=3$ is a constant. This situation is shown in Figure 3. The defender and pursuer are initially located $400$ m and $5000$ m radially apart from the evader, with respective LOS angles of $-45^{\circ}$ and $0^{\circ}$. The initial heading angles of the agents are $\gamma_{\mathrm{E}}=45^{\circ}$, $\gamma_{\mathrm{D}}=0^{\circ}$, and $\gamma_{\mathrm{P}}=180^{\circ}$. Under these settings, the initial value of the angle $\chi$ is $228^{\circ}$. It can be observed from Figure 3(a) that based on different values of the desired $\chi^{\star}$, the agents maneuver differently. However, the pursuer is always intercepted before it can capture the evader. Figure 3(b) depicts the sliding manifold and the error profiles of various desired angles $\chi^{\star}$, which evidences that $\underaccent{\bar}{\mathcal{S}}$ converges within 3 s and $\beta$ nullifies within 6 s, as expected. Therefore, regardless of the three-body initial engagement geometry, the evader-defender team maneuvers to ensure the pursuer’s capture. Figures 3(c) and 3(d) illustrate the various control efforts of the agents. It is important to note that the efforts of the evader and the defender (individually and jointly) have small magnitudes in the endgame.
Since the proposed defense strategy is independent of the pursuer’s maneuver, the defender is able to capture the pursuer even if it uses a different guidance law. Figure 4 portrays various cases when the pursuer uses proportional-navigation guidance (PNG), pure pursuit guidance (PPG), and deviated pursuit guidance (DPG) to capture the evader. The guidance command for PNG, PPG, and DPG are chosen to be $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)$, and $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}-\delta\right)$, respectively with $\mathscr{N}=3$ and $\delta=20^{\circ}$. For a fixed angle, $\chi^{\star}=180^{\circ}$, the defender always intercepts the pursuer midway and safeguards the evader. In Figure 4, the initial conditions are kept the same as that in the previous case, except $\chi^{\star}=180^{\circ}$. We also observe similar behaviors of the sliding manifold, error, and the agents’ control efforts.
We now consider cases when the evader is stationary, so the pursuer can head directly toward it with an optimal effort once any heading angle errors vanish. In Figure 5, the pursuer starts at a distance of 5000 m from the evader with a LOS angle of $2^{\circ}$. The defender, on the other hand, starts at three different positions in three different cases. With respect to the evader, the defender is at a relative separation of $400$ m, $1500$ m, and $2000$ m with LOS angles of $-45^{\circ}$, $-10^{\circ}$, and $10^{\circ}$, respectively. The defender aims to attain $\chi^{\star}=180^{\circ}$ in each case to arrive directly between the evader and the pursuer such that the latter’s capture can be guaranteed. We notice that even if the evader does not maneuver, the defender maneuvers accordingly to satisfy the specific geometrical conditions for a constant $\chi^{\star}$, thereby guaranteeing pursuit-evasion. On the other hand, the situation in Figure 6 assumes a fixed position of the defender ($r_{\mathrm{DE}}=400$ m, $\lambda_{\mathrm{DE}}=-45^{\circ}$), whereas the pursuer is located at different positions in different cases. The initial configurations of the pursuer relative to the evader are $r_{\mathrm{EP}}=5000$ m with LOS angles of $-5^{\circ},0^{\circ},$ and $2^{\circ}$, respectively. Once again, we observe that the fixed angle $\chi^{\star}$ is achieved by the defender in $t^{\star}=6$ s and maintained thereafter to ensure the pursuer’s capture. We now demonstrate the performance of the proposed strategy when the maximum steering controls of the defender and pursuer are chosen to be $8$ g, whereas the maximum steering control input for the evader is chosen to be $4$ g. While the other engagement settings are kept the same as for $\chi^{\star}=180^{\circ}$, and we depict the performance under reduced maximum capability through Figure 7(d). It can be observed that the defender intercepts the pursuer by following a similar behavior as when maximum accelerations are higher. This further bolsters the arguments presented in this work that pursuit-evasion is guaranteed regardless of the three-body initial engagement geometry and the knowledge of the pursuer’s strategy.
V Conclusions
In this paper, we have introduced a geometric approach that addresses the challenge of pursuit-evasion scenarios involving three agents with arbitrary initial geometries. Our proposed solution, the evader-defender cooperative guidance strategy, offers an effective means to guarantee pursuit-evasion under diverse conditions. Specifically, it guarantees that the defender reaches a specific angle precisely within a predefined time, irrespective of the initial engagement geometry, thus intercepting the pursuer before it can capture the evader. A distinguishing feature of our approach is its adaptability to scenarios characterized by nonlinear dynamics, non-holonomic constraints, and large heading angle errors, which are often encountered in practical motion control situations, including aircraft defense. This paper not only presents a robust and practical solution for guaranteed pursuit-evasion problems but also lays the foundation for identifying crucial conditions and configurations that facilitate successful evasion or capture. Analyzing geometrical solutions in three-dimensional settings could be an interesting future research.